# 3D-MolT5: Towards Unified 3D Molecule-Text Modeling with 3D Molecular Tokenization

### v1

@@ Abstract
The integration of molecule and language has garnered increasing attention in molecular science. Recent advancements in Language Models (LMs) have demonstrated potential for the comprehensive modeling of molecule and language. However, existing works exhibit notable limitations. Most existing works overlook the modeling of 3D information, which is crucial for understanding molecular structures and also functions. While some attempts have been made to leverage external structure encoding modules to inject the 3D molecular information into LMs, there exist obvious difficulties that hinder the integration of molecular structure and language text, such as modality alignment and separate tuning. To bridge this gap, we propose 3D-MolT5, a unified framework designed to model both 1D molecular sequence and 3D molecular structure. The key innovation lies in our methodology for mapping fine-grained 3D substructure representations (based on 3D molecular fingerprints) to a specialized 3D token vocabulary for 3D-MolT5. This 3D structure token vocabulary enables the seamless combination of 1D sequence and 3D structure representations in a tokenized format, allowing 3D-MolT5 to encode molecular sequence (SELFIES), molecular structure, and text sequences within a unified architecture. Alongside, we further introduce 1D and 3D joint pre-training to enhance the model’s comprehension of these diverse modalities in a joint representation space and better generalize to various tasks for our foundation model. Through instruction tuning on multiple downstream datasets, our proposed 3D-MolT5 shows superior performance than existing methods in molecular property prediction, molecule captioning, and text-based molecule generation tasks. Our code will be available on GitHub soon.
@@ 1 Introduction
Molecule plays a pivotal role in various scientific and industrial applications, ranging from pharmaceuticals to materials science [10, 9, 1]. In recent years, the development of Language Models (LMs) [4, 37, 52] has led to a growing focus on joint modeling of molecule and language [13, 61, 49, 25]. This is primarily driven by the extensive molecular-related textual descriptions available in biological literature [5, 57] and databases [19]. By training on these textual descriptions, LMs can capture comprehensive knowledge that aids in understanding molecules, thus better generalizing to molecule-related tasks, such as molecule-text retrieval [61, 49], molecule captioning, and description-guided molecule design [13, 29, 39]. Several studies have investigated molecule-text LMs along the way. Given that language itself represents a form of sequence, some attempts [13, 61, 39] employ autoregressive pre-training of LMs to simultaneously model 1D molecular sequences (e.g., SMILES [54], SELFIES [20, 21]) and text sequences. To further integrate 2D graph knowledge of molecule with text, some works leverage either contrastive pre-training between 2D molecular graphs and text [12, 49, 48, 33, 28], or align the 2D molecular graph encoders with LMs [30, 6] through multi-stage pre-training inspired by BLIP2 [24].
However, most existing works overlooked the molecular 3D structure, which contains critical stereochemical information, such as the spatial arrangement of atoms and functional groups. This information is crucial for tasks such as molecular energy prediction [47, 17], 3D function prediction [25], protein-ligand interaction [53], and so on. Recognizing this, several efforts have been made to integrate the 3D molecular information into LMs. The most representative ones, MolLM [51], 3D-MoLM [25], and MolBind [58], incorporate 3D molecular inputs with language by integrating external molecular structure encoders. They use external molecular encoders to first get the molecular representations, which are then aligned with LMs with continued training/tuning. Despite achieving preliminary success, these methods exhibit several shortcomings: (a) The molecular structure encoder is pre-trained separately, where the representation is not in the same space with LMs. (b) Though modality alignment can help mitigate the representation gap, it is challenging due to differences in model architectures, and accessing alignment quality is also difficult. (c) The final performance of the integrated model heavily relies on the performance of the external structure encoder, which can not be directly controlled.
@@
To address these limitations, we propose 3D-MolT5, a versatile T5 framework that can understand 3D Molecular structure and handle various 3D-dependent tasks with text instructions. To enable LMs to understand 3D molecular structure, we introduce a 3D molecular tokenization method based on the molecular E3FP (Extended 3D Fingerprint) [2] algorithm. E3FP can simply tokenize the 3D molecular structure into 3D tokens, with each token encapsulating the 3D information of a substructure centered around a specific atom. Additionally, since most 1D SELFIES tokens (e.g., [C] and [O]) represents specific atoms, the tokens from both 1D and 3D modalities can be directly aligned at the atomic level. Hence, the embeddings of the same atom in both 1D and 3D tokens are then summed to form the final joint representation, which can then enable the T5 model to effectively learn and comprehend molecular information by leveraging both the sequence and structure tokens, leading to improved performance across various molecular tasks.
@@
With tokenized 1D and 3D molecules, we conduct comprehensive molecule-text pre-training for our 3D-MolT5 framework. The pre-training tasks are inspired by the “T5 objective" from T5 [43] pre-training, which uses a “recover masked spans” objective that aims at recovering each masked span. In our 3D-MolT5, there are five types of pre-training tasks: 1) 1D denoising: apply T5 objective to SELFIES, text, and wrapped text, where molecules mentioned in the text are replaced with SELFIES. 2) 1D + 3D joint denoising: apply T5 objective to the summed 1D and 3D tokens, and the target is to recover the masked 1D SELFIES tokens. 3) 3D to 1D translation: given the 3D molecular tokens, the target is to generate the corresponding 1D SELFIES. 4) 3D molecule to text translation: given the summed 1D and 3D tokens, the target is to generate its textual description. 5) Text to 1D molecule translation: given the textual description, the target is to generate the corresponding 1D SELFIES.
@@
After pre-training, we conduct instruction tuning on various molecule-text tasks, including molecular property prediction (both 3D-dependent and 3D-independent) on PubChemQC [34, 59], QM9 [47, 14] and PubChem [19] datasets, molecule captioning (3D-dependent) on PubChem [19] dataset, and text-based molecule generation (3D-independent) on CheBI-20 dataset [13]. The results show that 3D-MolT5 achieves superior performance across these tasks, underscoring its versatility and efficacy in both 3D-dependent and 3D-independent molecule-text tasks.
@@ 2 Related Work
The 1D sequence is the most widely used form of molecular encoding, typically obtained by traversing the atoms in a molecular graph in a specified order. The simplified molecular-input line-entry system (SMILES) [54, 55] is the most used one, while Self-Referencing Embedded Strings (SELFIES) [20, 21] has recently gained popularity due to its strong robustness. 2D graph representations align naturally with molecular topological structures, as molecules inherently form 2D graphs with atoms serving as nodes and chemical bonds as edges. In contrast, 3D structures provide information about the spatial arrangement of atoms, offering valuable insights into molecular geometry and interactions. In addition to these representations, molecular fingerprints (FPs) are widely used, especially in molecular similarity searches and virtual screening [7]. FPs encode critical information about molecular structure as a sequence of binary bits, which are useful for property predictions [18, 56]. Common examples include Morgan FPs, such as extended-connectivity fingerprints (ECFPs) and functional class fingerprints (FCFPs) [45], as well as RDKit (topological) fingerprints [22]. However, these fingerprints primarily capture 2D topological features and do not account for 3D structural patterns. Spherical extended 3D fingerprints (E3FPs) [2] effectively incorporate neighboring atoms in 3D space to encode 3D information. Our structure-aware 3D molecular vocabulary is built on E3FP [2], which is then used for our atom-centric joint representation.
@@
Recent advances have integrated LMs with molecules to enhance the understanding of molecular structures and properties [62, 41]. MolT5 [13], BioT5 [39], and BioT5+ [40] are T5-based [43] models that jointly trained on 1D molecular sequences and text, then fine-tuned for molecule-related tasks. Mol-Instructions [14] and LlaSMol [60] provide instruction datasets with molecules represented as SMILES or SELFIES for instruction tuning. 2D molecular graphs are also used to inject topological knowledge to LMs through external graph encoding modules. For example, MoMu [49], MoleculeSTM [28], and MolFM [33] employ cross-modal contrastive learning on 2D molecular graphs and corresponding text. MolCA [30] aligns 2D molecular space with text space using cross-modal pre-training. Some recent works also consider 3D molecular information. For instance, MolBind [58] uses contrastive learning to align the 2D graph encoder, 3D structure encoder, and language encoder, showing strong performance in cross-modal retrieval tasks. MolLM [51] adopt a similar learning objective but use Transform-M-style [32] encoder to integrate both 2D and 3D molecular information. Following a similar paradigm to BLIP2 [24], 3D-MoLM [25] equips the LM with an external 3D molecular encoder and curates the 3D-MoIT dataset for 3D molecule-text instruction tuning. 3D-MoLM combines 1D SMILES and 3D molecular representation for 3D molecule-to-text interpretation. However, these methods do not model 1D molecular sequences, 3D molecular structures, and text in a unified manner, as the 3D molecule is encoded by an external module. This separation presents challenges in achieving a holistic integration of multimodal molecular information.
@@ 3 Methods
The overview of 3D-MolT5 is shown in Figure 1. We first give some preliminaries in Section 3.1, focusing on the 1D sequence representation of molecule. In Section 3.2, we present the 3D structure-aware E3FP fingerprint and how we adopt it in 3D-MolT5. We then introduce our 3D molecular tokenization and the integration with 1D tokenization in Section 3.3. Lastly, we present our multi-task pre-training framework in Section 3.4.
3.1 Preliminary
For a given molecule $M$, we use SELFIES [20] as its sequence representation, which is more robust than SMILES [54, 55]. In SELFIES, each token, typically representing an atom group (like [C] and [=N]) or structure directive (like [Ring1] and [=Branch1]), is enclosed within brackets and can be naturally tokenized based on these delimiters. Therefore, the 1D SELFIES sequence can be represented as $S=\{s_{i}\}_{i=1}^{m}$, where $s_{i}$ is the SELFIES token and $m$ is the length of sequence [^1] Typically, $m$ is large than the number of atoms $n$, as in SELFIES there are some structure directive tokens and other uninformative tokens.. To ensure that each molecule has a unique SELFIES, and consequently a unique atom flatten order, we employ the canonical form of SELFIES. We realize this by firstly getting the canonical SMILES using RDKit [22], then converting it to SELFIES.
@@ 3.1 Preliminary
Given a molecule $M$ with $n$ atoms, and one of its 3D structure $C$, we use the E3FP [2] algorithm to generate its 3D fingerprint $F$. $F$ represents a bit vector, with $|F|$ denoting its length. We flatten the atoms of $M$ into a canonical atom sequence $A=\{a_{i}\}_{i=1}^{n}$, where $a_{i}$ represents one of the heavy atoms. For each atom $a_{i}$, with $k$ as the number of iterations for the E3FP algorithm, we can derive its 3D token ${\bm{d}_{i}}$, which consists of $k+1$ nonnegative integers. This process, illustrated in Figure 2, includes several steps as follows. (1) Structure representation initialization. At iteration 0, we define a set of atomic invariants for each atom $a_{i}$ within $M$, including properties such as the number of heavy atom immediate neighbors, atomic number, mass, and charge. These atomic invariants are hashed into integers $\hat{d}_{i0}$ to establish unique identifiers for each atom $a_{i}$. (2) Iterative spherical shell expansion. In each iteration, we extend the radius of the spherical shells centered on each atom $a_{i}$ by a radius multiplier $r$, capturing all atoms within this expanded radius. The connectivity and spatial arrangement of these atoms are encoded relative to their central atom, incorporating both bonded and non-bonded interactions. This allows E3FP to capture the 3D structure of the molecule, including relative atomic orientations and distances that are not evident in simpler 2D representations. (3) Termination and outputs. Each iteration results in a new layer of structural information, extending until either all atoms in $M$ are included or a predefined maximum number of iterations $k$ is reached. For each atom $a_{i}$, we combine its substructure information in each layer by concatenating the hashed integers, resulting in the vector $\hat{\bm{d}_{i}}=[\hat{d}_{i0},\hat{d}_{i1},\dots,\hat{d}_{ik}]$. Each item of $\hat{\bm{d}_{i}}$ is then converted to $\bm{d}_{i}$ by taking its value modulo $|F|$, expressed mathematically as $\bm{d}_{i}=\hat{\bm{d}_{i}}\mod|F|$. In 3D-MolT5, we do not use the E3FP $F$ but instead use the 3D tokens $\bm{d}_{i}$ for each atom $a_{i}$, which are then integrated with 1D sequence. Specifically, the $k$ is set to 3, as our preliminary experiments indicate that three iterations are sufficient for the E3FP algorithm to converge, capturing all potentially occurring substructures for the vast majority of molecules. The $|F|$ is set to 4096 and $r$ is set to 1.718 Å. For more details of the E3FP algorithm, we suggest to refer the original paper [2].

### v2

# 3D-MolT5: Leveraging Discrete Structural Information for Molecule-Text Modeling

@@ Abstract
The integration of molecular and natural language representations has emerged as a focal point in molecular science, with recent advancements in Language Models (LMs) demonstrating significant potential for comprehensive modeling of both domains. However, existing approaches face notable limitations, particularly in their neglect of three-dimensional (3D) information, which is crucial for understanding molecular structures and functions. While some efforts have been made to incorporate 3D molecular information into LMs using external structure encoding modules, significant difficulties remain, such as insufficient interaction across modalities in pre-training and challenges in modality alignment. To address the limitations, we propose 3D-MolT5, a unified framework designed to model molecule in both sequence and 3D structure spaces. The key innovation of our approach lies in mapping fine-grained 3D substructure representations into a specialized 3D token vocabulary. This methodology facilitates the seamless integration of sequence and structure representations in a tokenized format, enabling 3D-MolT5 to encode molecular sequences, molecular structures, and text sequences within a unified architecture. Leveraging this tokenized input strategy, we build a foundation model that unifies the sequence and structure data formats. We then conduct joint pre-training with multi-task objectives to enhance the model’s comprehension of these diverse modalities within a shared representation space. Thus, our approach significantly improves cross-modal interaction and alignment, addressing key challenges in previous work. Further instruction tuning demonstrated that our 3D-MolT5 has strong generalization ability and surpasses existing methods with superior performance in multiple downstream tasks, such as nearly 70% improvement on the molecular property prediction task compared to state-of-the-art methods. Our code is available at https://github.com/QizhiPei/3D-MolT5.
@@ 1 Introduction
Molecule plays a pivotal role in various scientific and industrial applications, spanning from pharmaceuticals to materials science (Drews, 2000; Dara et al., 2022; AI4Science & Quantum, 2023). In recent years, the development of Language Models (LMs) (Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024) has garnered significant attention towards the joint modeling of molecule and language (Edwards et al., 2022; Zeng et al., 2022; Li et al., 2023c). LMs trained on textual descriptions of molecules can acquire comprehensive knowledge that enhances molecular understanding, thereby improving generalization to various molecule-related tasks, such as molecule-text retrieval (Zeng et al., 2022; Su et al., 2022) and molecule captioning (Edwards et al., 2022; Liu et al., 2023b; Pei et al., 2023). Language, inherently sequential, has inspired researchers to explore autoregressive pre-training of LMs for jointly modeling molecular sequences (e.g., SMILES (Weininger, 1988), SELFIES (Krenn et al., 2020; 2022)) and text sequences (Edwards et al., 2022; Zeng et al., 2022; Pei et al., 2023). To incorporate 2D graph information, two primary approaches have emerged: contrastive pre-training between 2D molecular graphs and text (Edwards et al., 2021; Su et al., 2022; Seidl et al., 2023; Luo et al., 2023; Liu et al., 2023a), and alignment of 2D molecular graph encoders with LMs (Liu et al., 2023c; Cao et al., 2023) through multi-stage pre-training inspired by BLIP2 (Li et al., 2023b).
However, most existing works have overlooked the molecular 3D structure, which contains crucial stereochemical information for function-related tasks (Ruddigkeit et al., 2012; Hu et al., 2021; Li et al., 2023c; Wang et al., 2005). Few attempts (Tang et al., 2023; Li et al., 2023c; Xiao et al., 2024; Zhao et al., 2024) try to eliminate this limitation by integrating external molecular structure encoders to incorporate the 3D molecular inputs with language. Through alignment training between the external molecular structure encoders and LMs, these approaches achieve preliminary success, but notable shortcomings exist in these methods: (1) Insufficient interaction across modalities in pre-training: The molecular structure encoder is pre-trained separately from the text, resulting in inadequate interaction between different modalities during pre-training. (2) Challenges in modality alignment: The different representation spaces of pre-trained molecular encoders and LMs require alignment training. However, modality alignment is always challenging (Baltrušaitis et al., 2018) for different reasons (e.g., limited molecule-text paired data, discrete text tokens, and continuous 3D structure representation). Besides, how to assess the quality of alignment well remains unclear. (3) Dependency on external encoder: Though it is efficient to incorporate the pre-trained external structure encoder, the performance of the encoder can not be directly controlled within the framework, which also raises difficulty in performance alignment.
@@
To address these limitations, inspired by the joint multi-modal modeling in Vision-Language (Team, 2024; Xie et al., 2024; Zhou et al., 2024), we propose 3D-MolT5, a versatile T5 framework capable of understanding 3D Molecular structure to handle various 3D-dependent tasks simultaneously with text instructions. To enable LMs to comprehend 3D molecular structures, the crucial innovation is that we introduce a 3D molecular tokenization method based on the Extended 3D Fingerprint (E3FP) algorithm (Axen et al., 2017). Specifically, E3FP tokenizes the 3D molecular structure into discrete 3D tokens, with each token encapsulating the 3D information of a substructure centered around a specific atom. Since most 1D SELFIES tokens (e.g., [C] and [O]) represent specific atoms, the tokens from both 1D and 3D modalities can be directly aligned at the atomic level. The embeddings of the same atom in both 1D and 3D tokens are then summed to form the final joint representation. In this way, it enables effective learning of molecular information by leveraging both sequence and structure tokens for molecules, allowing the representation of 1D molecule, 3D molecule, and 1D text modalities all using discrete tokens, such that all modalities can be easily trained in LMs. Therefore, we not only remove the dependency and requirement on external molecular structure encoders but also eliminate the necessity of challenging modality alignment training.
@@
With tokenized 1D and 3D molecules, we conduct comprehensive molecule-text pre-training on our 3D-MolT5 framework. The pre-training tasks are inspired by the “T5 objective” (Raffel et al., 2020), which employs a “recover masked spans” objective. In 3D-MolT5, we design five types of pre-training tasks: (1) 1D denoising: Apply T5 objective to SELFIES, text, and wrapped text, where molecules mentioned in the text are replaced with SELFIES. (2) 1D + 3D joint denoising: Apply T5 objective to the summed 1D and 3D tokens, with the target being to recover the masked 1D SELFIES tokens. (3) 3D to 1D translation: Given the 3D molecular tokens, generate the corresponding 1D SELFIES. (4) 3D molecule to text translation: Given the summed 1D and 3D tokens, generate its textual description. (5) Text to 1D molecule translation: Given the textual description, generate the corresponding 1D SELFIES. Consequently, our 3D-MolT5 pre-training allows for early and extensive interaction between different modalities in the pre-training stage, enhances representation, and better integrates information across modalities.
@@
To verify our 3D-MolT5 framework, we conduct instruction tuning after pre-training on various molecule-text tasks, including molecular property prediction (both 3D-dependent and 3D-independent), molecule captioning (3D-dependent), and text-based molecule generation (3D-independent). The results show that both the Specialist (single-task tuned) and Generalist (multi-task tuned) versions of 3D-MolT5 achieve superior performance across these tasks. For example, on the 3D-dependent molecular property prediction task with the PubChemQC (Maho, 2015) dataset, 3D-MolT5 achieves an improvement of nearly 70% compared to 3D-MoLM (Li et al., 2023c). These results underscore the versatility and efficacy of our 3D-MolT5 in both 3D-dependent and 3D-independent molecule-text tasks.
@@ 2 Related Work
The 1D sequence is the most widely used form of molecular encoding, typically obtained by traversing the atoms in a molecular graph in a specified order. The simplified molecular-input line-entry system (SMILES) (Weininger, 1988; Weininger et al., 1989) is the most common, while Self-Referencing Embedded Strings (SELFIES) (Krenn et al., 2020; 2022) has recently gained popularity due to its robust nature. 2D graph representations align naturally with molecular topological structures, as molecules inherently form 2D graphs with atoms serving as nodes and chemical bonds as edges (Guo et al., 2023). In contrast, 3D structures provide information about the spatial arrangement of atoms, offering valuable insights into molecular geometry and interactions. Molecular fingerprints (FPs) are also widely used, especially in molecular similarity searches and virtual screening (Cereto-Massagué et al., 2015). FPs encode critical information about molecular structure as a sequence of binary bits, which are useful for property predictions (Jeon & Kim, 2019; Wen et al., 2022). Common examples include Morgan FPs, such as extended-connectivity fingerprints (ECFPs) and functional class fingerprints (FCFPs) (Rogers & Hahn, 2010a), as well as RDKit (topological) fingerprints (Landrum et al., 2023). However, these fingerprints primarily capture 2D topological features and do not account for 3D structural patterns. Spherical extended 3D fingerprints (E3FPs) (Axen et al., 2017) effectively incorporate neighboring atoms in 3D space to encode 3D information. Our structure-aware 3D molecular vocabulary is built on E3FP (Axen et al., 2017), which is then used for our atom-centric joint representation.
@@
Recent advancements have integrated LMs with molecules to enhance the understanding of molecular structures and properties (Zhang et al., 2024b; Pei et al., 2024b; Taylor et al., 2022). MolT5 (Edwards et al., 2022), BioT5 (Pei et al., 2023), and BioT5+ (Pei et al., 2024a) are T5-based (Raffel et al., 2020) models that jointly trained on 1D molecular sequences and text sequences, followed by fine-tuning for tasks related to molecules. Mol-Instructions (Fang et al., 2023) and LlaSMol (Yu et al., 2024) offer instruction datasets where molecules are represented as SMILES or SELFIES for instruction tuning. Additionally, 2D molecular graphs have been utilized to infuse topological knowledge into LMs via external graph encoding modules. For instance, MoMu (Su et al., 2022), MoleculeSTM (Liu et al., 2023a), and MolFM (Luo et al., 2023) employ cross-modal contrastive learning on 2D molecular graphs and corresponding text. MolCA (Liu et al., 2023c) and MolX (Le et al., 2024) align 2D molecular space with text space through cross-modal pre-training. UniMoT (Zhang et al., 2024a) proposes a Vector Quantization-driven tokenizer to convert the 2D molecular graphs into molecule tokens, aiming at unified modeling of molecule and text. Recent endeavors have also incorporated 3D molecular information. MolBind (Xiao et al., 2024), for example, employs contrastive learning to align the 2D graph encoder, 3D structure encoder, and language encoder, demonstrating strong performance in cross-modal retrieval tasks. In line with the BLIP2 (Li et al., 2023b) paradigm, 3D-MoLM (Li et al., 2023c) equips the LM with an external Uni-Mol encoder (Zhou et al., 2023) and curates the 3D-MoIT dataset for 3D molecule-text instruction tuning. 3D-MoLM combines 1D SMILES and 3D molecular representations for 3D molecule-to-text interpretation. Nonetheless, these methods do not unify the modeling of molecular sequences, molecular structures, and text sequences, as the 3D molecules are encoded by an external module, posing challenges in attaining a comprehensive integration of multimodal molecular information.
@@ 3 Methods
The overview of 3D-MolT5 is shown in Figure 1. We first introduce the sequence representation of molecules in Section 3.1, as it is the preliminary knowledge for our molecular tokenization. In Section 3.2, we present the 3D structure-aware E3FP fingerprint and how we adopt it in 3D-MolT5. We then introduce our 3D molecular tokenization and the integration with 1D tokenization in Section 3.3. Lastly, we present our multi-task pre-training framework in Section 3.4.
3.1 1D Sequence Representation
The 1D sequence representation for molecules lays the groundwork for our molecular tokenization (Section 3.3), hence we give the necessary descriptions here. In this work, for a given molecule $M$, we use SELFIES (Krenn et al., 2020) as its sequence representation, offering enhanced robustness and validity compared to SMILES (Weininger, 1988; Weininger et al., 1989). In SELFIES, each token, generally denoting an atom group (like [C] and [=N]) or structure directive (like [Ring1] and [=Branch1]), is enclosed within brackets, facilitating straightforward tokenization based on these demarcations. Therefore, the 1D SELFIES sequence can be represented as $S=\{s_{i}\}_{i=0}^{m-1}$, where $s_{i}$ is a SELFIES token and $m$ denotes sequence length [^1] Typically, $m$ is large than the number of atoms $n$, as in SELFIES there are some structure directive tokens.. To ensure that each molecule has a unique SELFIES, and thus a unique atom flatten order, we employ the canonical form of SELFIES.
@@
In our 3D tokenization, the focus is on transforming the continuous 3D molecular structure into discrete tokens. To achieve this, we leverage the 3D molecular fingerprint E3FP (Axen et al., 2017), which efficiently converts 3D structures into discrete identifiers, enabling a tokenized representation of spatial molecular information. For a molecule $M$ composed of $n$ atoms and one of its 3D conformers, the E3FP algorithm generates a 3D fingerprint $F$, represented as a bit vector, with $|F|$ indicating its length. We transform the atoms of $M$ into a canonical atom sequence $A=\{a_{i}\}_{i=0}^{n-1}$, where $a_{i}$ represents one of the heavy atoms. For each atom $a_{i}$, with $k$ representing the number of iterations for the E3FP algorithm, we can derive its 3D token ${\bm{d}_{i}}$, composed of $k+1$ non-negative identifiers. This process, illustrated in Figure 2 and Algorithm 1 (some details and special cases are omitted and provided in the Appendix A), encompasses several steps as follows. (1) Structure representation initialization. At iteration 0, we establish a set of atomic invariants $A_{i}$ for each atom $a_{i}$ within $M$, including attributes such as atomic number, the number of immediate neighbors, and whether the atom is part of a ring, among others. These atomic invariants are hashed into identifier $\hat{d}_{i,0}$ by MurmurHash3 (Appleby, 2016) algorithm to create unique identifiers for each atom $a_{i}$. (2) Iterative spherical shell expansion. In each iteration $j$, we expand the radius of the spherical shells centered on each atom $a_{i}$ by a radius multiplier $r$, including all atoms within the expanded radius $R=r\cdot j$. The connectivity identifier, $c_{k}^{i}$, derived from Connectivity, and the stereochemical identifier, $s_{k}^{i}$, obtained from Stereochemistry, of these neighboring atoms are encoded relative to the central atom, incorporating both bonded and non-bonded interactions. These identifiers allow E3FP to capture the 3D molecular structure, including relative atomic orientations and distances, which are absent in 2D representations. A detailed explanation of Connectivity and Stereochemistry is provided in Appendix A.1. The current iteration number $j$, the identifier of $a_{i}$ from the previous iteration, and the neighbors’ information ($c_{k}^{i}$, $\hat{d}_{k,j-1}$ $s_{k}^{i}$) are combined and hashed into the identifier $\hat{d}_{i,j}$. Each iteration produces a new layer of structural information, continuing until either a predefined maximum number of iterations $k$ is reached or all atoms in $M$ are included in the current shell.
Algorithm 1 E3FP Algorithm. $\cup$ represents the concatenation operation.
1:  Input: Molecule $M$, maximum iteration number $k$, shell radius multiplier $r$
2:  Initialize $\mathbf{D}\leftarrow\{\}$
3:  // Step 1: Structure Representation Initialization
4:  for each atom $a_{i}$ in $M$ do
5:     Initialize atomic invariants $A_{i}$
6:     $\hat{d}_{i,0}\leftarrow\textit{MurmurHash3}(A_{i})$
7:     $\bm{\hat{d}_{i}}\leftarrow[\hat{d}_{i,0}]$
8:  end for
9:  // Step 2: Iterative Spherical Shell Expansion
10:  for iteration $j=1$ to $k$ do
11:     $R\leftarrow r\cdot j$
12:     for each atom $a_{i}$ in $M$ do
13:        $L\leftarrow[j,\hat{d}_{i,j-1}]$
14:        for each neighbor $a_{k}$ within radius $R$ do
15:           $c_{k}^{i}\leftarrow\textit{Connectivity}(a_{k},a_{i})$
16:           $s_{k}^{i}\leftarrow\textit{Stereochemistry}(a_{k},a_{i})$
17:           $L\leftarrow L\cup[c_{k}^{i},\hat{d}_{k,j-1},s_{k}^{i}]$
18:        end for
19:        $\hat{d}_{i,j}\leftarrow\textit{MurmurHash3}(L)$
20:        $\bm{\hat{d}_{i}}\leftarrow\bm{\hat{d}_{i}}\cup[\hat{d}_{i,j}]$
21:     end for
22:  end for
23:  // Step 3: Folding
24:  for each atom $a_{i}$ in $M$ do
25:     $\bm{d_{i}}\leftarrow\bm{\hat{d}_{i}}\mod|F|$
26:     $\mathbf{D}\leftarrow\mathbf{D}\cup d_{i}$
27:  end for
28:  Output: 3D structure identifier matrix $\mathbf{D}$
(3) Folding. For each atom $a_{i}$, we aggregate its substructure information from each layer by concatenating the hashed identifiers, resulting in the vector $\hat{\bm{d}_{i}}=[\hat{d}_{i,0},\hat{d}_{i,1},\dots,\hat{d}_{i,k}]$. Each 32-bit element of $\hat{\bm{d}_{i}}$ is then “folded” down to $|F|$ bit by applying the modulo operation, mathematically represented as $\bm{d}_{i}=\hat{\bm{d}_{i}}\mod|F|$. In 3D-MolT5, we do not use the E3FP $F$ directly but instead employ the 3D tokens $\bm{d}_{i}$ for each atom $a_{i}$, which are then integrated with the 1D sequence. More details and analysis about the E3FP algorithm, including connectivity and stereochemistry encoding, SE(3)-invariance, time and space complexity, the hyperparameter settings, special cases, a specific example for better illustration, information loss brought by discrete representation, are introduced in Appendix A.